OpenStudy (anonymous):

find the average value of f(x) = x^2 on the interval [1, 6]

OpenStudy (anonymous):

\[\frac{f(6) -f(1)}{6-1}\]

OpenStudy (anonymous):

is it as simple as that? i thought we might have to integrate

OpenStudy (anonymous):

how do i do that, please solve it out

OpenStudy (anonymous):

surely that answer is finding the average value of a line through 6,f(6) and 1,f(1)

OpenStudy (anonymous):

The formula he used is for the mean value theorem, the average value though would be an integral I believe.

OpenStudy (anonymous):

Oh sorry, I found the average rate of change.. My mistake!

OpenStudy (anonymous):

In this case we need to use: \[f_{avg} = \frac{1}{6-1}\int\limits_{1}^{6} x^2 dx\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You are ok from there?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Whats the problem?

OpenStudy (anonymous):

i keep getting the wrong answer, i thought it was 7 but its not

OpenStudy (anonymous):

Post your steps.

OpenStudy (anonymous):

i wrote them out on paper, its fine really

OpenStudy (anonymous):

\[\frac{1}{5}*(\frac{6^3}{3}-\frac{1^3}{3})\]